a circuit consist of an inductance of 0.05H, a resistance of 20ohm, a capacitor of 100uF and emf of 100cos200t.find (i) and (q) given that q=0and i=0at t=0

Grade:

1 Answers

Jit Mitra
12 years ago

Z = sqrt(20^2 + (200*0.05 - 1/(200*100*10^-6))^2) = sqrt(400 + (10-50)^2) = sqrt(2000) = 20√5

 

I = E/Z = √5 * cos(200t)

 

q = (integration from t=0 to t=t) I.dt = √5/200 * [sin(200t)]